### v1

# Nucleon Consumption and Mass-Energy Conversion Induced by Dark Matter

Abstract
We propose the nucleon consumption induced by dark matter as a new scenario to overcome the energy threshold of direct detection. It can be realized with proton ($\chi+p\rightarrow\chi+\ell^{+}$) or neutron ($\chi+n\rightarrow\chi+\nu$) target. Both effective operators and concrete models are provided to illustrate the idea. Since the initial DM and nucleon velocity is only $10^{-3}$ and $1/4$ of the speed of light, respectively, the cross section and hence event rate are determined by the involved particle masses and not affected by the nucleon Fermi motion. Of the two realizations, the proton consumption has richer phenomena with both charged lepton and the daughter nuclei de-excitation to allow double or even triple coincidence to significantly suppress the background. We also illustrate the projected sensitivities at DM and neutrino experiments such as PandaX-4T, DUNE, JUNO, and Hyper-K.
Introduction – The nature of dark matter (DM) is one of the biggest unknowns in our Universe Young:2016ala ; Arbey:2021gdg ; Fairbairn:2022gar . The particle DM is a well-established paradigm that can explain the observed astrophysical and cosmological phenomena with minimal parameter set. Especially, the weakly interacting massive particle (WIMP), whose stability is usually protected by discrete symmetry, is a theoretically well-motivated scenario that can meet the required properties to address the DM conundrum and at the same time have a detectable possibility in terrestrial experiments Roszkowski:2017nbc ; Schumann:2019eaa . However, the DM direct detection experiments have not found signal but constrain the DM-nucleus cross section to an unprecedented level for the WIMP mass from GeV to hundreds of GeV Billard:2021uyg ; Misiaszek:2023sxe .
Since the DM-nucleus elastic scattering can only convert the DM kinetic energy to the nucleus recoil energy, it is difficult for a sub-GeV light DM to overcome the energy threshold of a direct detection experiment. To solve this issue, novel low-threshold experiments and theoretical ideas are proposed. On the experimental side Essig:2022dfa ; Mitridate:2022tnv , for instance, the DM-electron scattering Essig:2011nj ; Essig:2012yx , the Migdal effect Ibe:2017yqa , and the bremsstrahlung processes Kouvaris:2016afs can enhance the signal strength. On the theoretical side, various DM scenarios have appeared.
The recently proposed fermionic absorption DM is another example of light DM particle that can inject enough energy to the detector to produce a detectable signal with either nuclei Dror:2019onn ; Dror:2019dib ; Li:2022kca ; Ma:2024aoc ; Ge:2024euk ; Ma:2024tkt or electron Dror:2020czw ; Ge:2022ius ; Li:2022kca ; Ge:2023wye target. The idea is that the DM ($\chi$) is absorbed by nuclear or atomic target ($T$) while emitting an undetectable neutrino ($\nu$), $\chi+T\to\nu+T^{\prime}$, while the target ($T^{\prime}$) recoils. Such signals have been searched at PandaX-4T PandaX:2022osq ; PandaX:2022ood , Majorana Majorana:2022gtu , EXO-200 EXO-200:2022adi , and CDEX CDEX:2022rxz ; CDEX:2024bum .
Using the fermion absorption and converting its mass to recoil energy for detecting an unkown particle has a long history back to the 1940s. Although Pauli also pointed out the possibility of using its electromagnetic dipole interactions to probe the neutrino he proposed in the same letter Pauli:1991owm , both its very existence and electromagnetic interactions are hypothetical. With only weak interactions, neutrinos are extremely difficult to detect. So Kan Chang Wang proposed in 1942 a very smart way of using the $K$-shell electron capture and detecting the mono-energetic daughter nuclei recoil to probe the existence of neutrino Wang:1942huv ; Wang:1947 . This is essentially absorbing/consuming the known electron and releasing its mass as energy to probe the unknown neutrino.
In this paper, we propose a new absorption mechanism by consuming nucleon rather than the DM particle, to overcome the detection threshold. Unlike the fermionic absorption DM, which is unstable and suffers from stringent indirect detection constraints due to decaying into SM light particles like neutrinos or photons (and possibly electron-positron pair or light mesons were these channels kinematically available) Dror:2019onn ; Dror:2019dib ; Dror:2020czw ; Ge:2022ius ; Ge:2023wye , the nucleon consuming DM is stable due to a $\mathbb{Z}_{2}$ symmetry protecting it from decaying. Since one nucleon is consumed in the process to release almost 1 GeV energy, we anticipate there will be enough energy deposit in the detector and thus can be searched for at the DM direct detection and neutrino experiments.
Nucleon Consumption Operators and Models – For the nucleon consumption scenarios induced by DM, it may happen at the nucleus level if the momentum transfer is not large enough. Then the target nucleus ($N$) reduces to a lighter one ($N^{\prime}$), $\chi+N\rightarrow\chi+N^{\prime}+\ell/\nu$ where the final-state charged lepton ($\ell$) or neutrino ($\nu$) is necessary to balance the fermion flow and angular momentum. However, it is not easy to make the momentum transfer transparent with a three-body final state. On the other hand, a large enough momentum transfer would render the nucleon consumption process to happen at the nucleon level. Since the DM particle remains intact, the consumed nucleon in the initial state will become either a positively charged lepton ($\chi+p\to\chi+\ell^{+}$) or a neutrino ($\chi+n\to\chi+\nu$). For both cases, the massive nucleon (either proton $p$ or neutron $n$) will release its mass around 938 MeV to a light charged lepton (positron $e^{+}$ or postively charged muon $\mu^{+}$) whose mass is much smaller or even a massless neutrino. With a two-body final state, the momentum transfer $q^{2}\equiv(p_{N}-p_{\ell})^{2}=-m_{\chi}(m^{2}_{N}-m^{2}_{\ell})/(m_{N}+m_{\chi})$ is fixed by the DM ($m_{\chi}$), nucleon ($m_{N}$), and lepton ($m_{\ell}$ for charged lepton or neutrino) masses. With both $m_{\chi}$ and $m_{N}$ around the GeV scale, the momentum transfer can probe a length scale much smaller than the nuclei size. In other words, the consumption scenario really happens at the nucleon level rather than with the nucleus as a whole.
For the DM mass below half of the nucleon mass, the same interaction that induces nucleon consumption can also lead to the proton ($p\rightarrow 2\chi+\ell^{+}$) or neutron ($n\to 2\chi+\nu$) decay. The subsequent dark decay of neutron can address the longstanding discrepancy between the bottle- and beam-type measurements of the neutron lifetime ParticleDataGroup:2022pth ; Yue:2013qrc . However, the nucleon decay can be forbidden by requiring a heavier DM mass to be on the safe side.
With a nonrelativistic incoming DM particle and the initial proton at rest, the consumed proton mass is released into the final-state DM and charged lepton energies $E_{\chi}=(m^{2}_{p}+2m_{p}m_{\chi}+2m^{2}_{\chi}-m^{2}_{\ell})/2(m_{p}+m_{\chi})$ and $E_{\ell}=(m^{2}_{p}+2m_{p}m_{\chi}+m^{2}_{\ell})/2(m_{p}+m_{\chi})$, respectively. Both energies are fully determined by the proton ($m_{p}$), DM ($m_{\chi}$), and lepton ($m_{\ell}$) masses. Although the muon mass is not that small, its value is just around 106 MeV and can hence be safely ignored as a good enough approximation. The proton realization can then be detected by the charged lepton and de-exicitation of the daughter nucleus.
While the charged lepton $e^{+}$ or $\mu^{+}$ cannot be replaced by another charged particle, the SM neutrino produced by $\chi+n\to\chi+\nu$ will escape detection and can be faked by any other neutral particles such as the sterile neutrino whose mass is also small enough or a light dark particle. Of the two realizations, the neutron consumption with only the daughter nucleus de-excitation leaves less signal in the detector. So we would focus on the proton consumption scenario for illustration while most results also apply for the neutron case.
The nucleon consumption can be described in terms of four-fermion effective operators involving a pair of fermion DM particles, proton, and charged lepton are,
${\cal O}_{\chi\ell p}^{\tt SS(P)}$ $=\overline{\chi}\chi\,\overline{\ell^{c}}(i\gamma_{5})p,$ (1a)
${\cal O}_{\chi\ell p}^{\tt PS(P)}$ $=\overline{\chi}i\gamma_{5}\chi\,\overline{\ell^{c}}(i\gamma_{5})p,$ (1b)
${\cal O}_{\chi\ell p}^{\tt VV(A)}$ $=\overline{\chi}\gamma_{\mu}\chi\,\overline{\ell^{c}}\gamma^{\mu}(\gamma_{5})p,$ (1c)
${\cal O}_{\chi\ell p}^{\tt AV(A)}$ $=\overline{\chi}\gamma_{\mu}\gamma_{5}\chi\,\overline{\ell^{c}}\gamma^{\mu}(\gamma_{5})p,$ (1d)
${\cal O}_{\chi\ell p}^{\tt TT(\tilde{T})}$ $=\overline{\chi}\sigma_{\mu\nu}\chi\,\overline{\ell^{c}}\sigma^{\mu\nu}(i\gamma_{5})p,$ (1e)
where $\ell^{c}$ is the charge conjugation of the lepton field. For the superscript, the first character stands for the Lorentz type of the DM current while the second and the one in parenthesis for the proton-lepton current. The DM current can also take the fermion number violating case, $\overline{\chi^{c}}\Gamma\chi$. Then only the scalar, pseudo-scalar, and axial-vector operators ${\cal O}_{\chi\ell p}^{\tt S,P,A}$ can survive. The strength of each operator is parametrized by a Wilson coefficient $C_{\chi\ell p}^{i}$. For the DM-neutron consumption, one may just replace the charged lepton-proton current by the neutrino-neutron current, $(\overline{\nu}\Gamma n)$ or $\overline{\nu^{c}}\Gamma n$.
The nucleon consumption operators in Eq. (1) can be realized by extending the SM with a Dirac fermion DM candidate $\chi(1,1,0)$ besides three leptoquarks $S_{1}(3,1,2/3)$, $S_{2}(3,1,-1/3)$, and $S_{3}(3,1,-1/3)$. In addition, one needs to impose a discrete $\mathbb{Z}_{4}$ symmetry under which the charge assignments are,
$\chi:i,\quad S_{1,2}:-i,\quad S_{3},L,e:-1,$ (2)
while the remaining SM particles are $\mathbb{Z}_{4}$ singlets. Here $L$ and $e$ are the SM left- and right-handed leptons. The relevant new physics Lagrangian is
${\cal L}_{\tt NP}$ $\ni y_{1}(\overline{u}\chi_{L})S_{1}+y_{2}(\overline{d}\chi_{L})S_{2}+y_{3}(\overline{Q}\epsilon L^{c})S_{3}$
$+y_{4}(\overline{u}e^{c})S_{3}-\mu_{S}\epsilon_{ijk}S_{1}^{i}S_{2}^{j}S_{3}^{k}+{\rm H.c.},$ (3)
where $Q,u,d$ are the SM quarks while $i,j,k$ are the color indices in the fundamental representation of $\rm SU(3)_{c}$. For the Yukawa couples $y_{1,2,3,4}$, the flavor indices have been omitted for simplicity. The dimensional $\mu_{S}$ is the triple scalar coupling.
[…]
Figure 1: The Feynman diagram for generating the effective operators responsible for the nucleon consumption induced by DM.
The Feynman diagram responsible for the nucleon comsuming DM is shown in Fig. 1, which will generate operators with (pseudo-)scalar DM currents in Eq. (1) once quarks condensate into nucleons. While the couplings with $S_{1}$ and $S_{2}$ already contribute one up and one down quarks, whether the model induces a proton or neutron consumption is determined by the $S_{3}$ Yukawa couplings. A vanishing $y_{3}$ prevents neutrino from participating the process and hence only leads to the proton consumption scenario with nonzero $y_{4}$. On the other hand, a nonzero $y_{3}$ can induce both proton and neutron consumption. Cancellation between the $y_{3}$ and $y_{4}$ contributions to the proton decay can also happen by tuning values. However, it is more desirable to adjust the model with a light sterile neutrino $\nu_{R}(1,1,0)$ that has a $\mathbb{Z}_{4}$ charge $-1$. Furthermore, the SM leptons and quarks all are neutral under the $\mathbb{Z}_{4}$ symmetry. In this case, the Yukawa terms associated with $y_{3,4}$ are replaced by $y_{3}(\overline{d}\nu_{R}^{c})S_{3}$.
Consumption at Rest and Fermi Motion – Since the halo DM particles travel at only $10^{-3}$ of the speed of light, it is a good approximation to treat them as at rest. The proton consumption process then happens at rest which leads to mono-energetic DM and charged lepton in the final state. Reparametrizing the DM mass as $x\equiv m_{\chi}/m_{p}$, the consumption cross section takes the form as,
$\sigma^{\tt SS(P)}_{\chi\ell p}v_{\chi}$ $\approx\frac{m^{2}_{p}}{32\pi\Lambda^{4}_{\tt eff}}\frac{(1+2x)^{4}}{(1+x)^{4}},$ (4a)
$\sigma^{\tt PS(P)}_{\chi\ell p}v_{\chi}$ $\approx\frac{m^{2}_{p}}{32\pi\Lambda^{4}_{\tt eff}}\frac{(1+2x)^{2}}{(1+x)^{4}},$ (4b)
$\sigma^{\tt VV(A)}_{\chi\ell p}v_{\chi}$ $\approx\frac{m^{2}_{p}}{16\pi\Lambda^{4}_{\tt eff}}\frac{(1+2x)^{2}(3+4x+2x^{2})}{(1+x)^{4}},$ (4c)
$\sigma^{\tt AV(A)}_{\chi\ell p}v_{\chi}$ $\approx\frac{m^{2}_{p}}{16\pi\Lambda^{4}_{\tt eff}}\frac{(1+2x)^{2}(3+8x+6x^{2})}{(1+x)^{4}},$ (4d)
$\sigma^{\tt TT(\tilde{T})}_{\chi\ell p}v_{\chi}$ $\approx\frac{m^{2}_{p}}{4\pi\Lambda^{4}_{\tt eff}}\frac{(1+2x)^{2}(5+10x+6x^{2})}{(1+x)^{4}},$ (4e)
where the Wilson coefficients $C_{\chi ep}^{i}\equiv\Lambda_{\tt eff}^{-2}$ can be expressed in terms of the effective scale $\Lambda_{\tt eff}$. For convenience, one may factorize out the $x$-dependent part, $\sigma^{i}_{\chi\ell p}v_{\chi}\equiv f^{i}(x)\times m^{2}_{p}/16\pi\Lambda^{4}_{\tt eff}$. As shown in Fig. 2, the $x$-dependent $f(x)/x$ that finally appears in the event rate is typically in the range of $10^{-3}\sim 10$. The size of $f(x)/x$ is sizable with the only exception of the pseudo-scalar case. The typical size of the consumption cross section, $\sigma_{0}\equiv m^{2}_{p}/16\pi\Lambda^{4}_{\tt eff}$, is mainly determined by the proton mass $m_{p}$ and the effective scale $\Lambda_{\tt eff}$.
Figure 2: The dependence of $f(x)/x$ as a function of $x=m_{\chi}/m_{p}$. Its value is sizable with the only exception of the pseudo-scalar case.
Since the consumption process happens at the nucleon level, the Fermi motion of the bound nucleon should be taken into consideration. The nucleon momentum distribution in nuclei can be described by the Fermi-gas model together with a high-momentum tail accounting for the nucleon-nucleon correlations Bodek:1980ar ,
$\hat{\rho}(|\boldsymbol{p}|)=\frac{1}{C}\begin{cases}1-6\left(k_{F}a\over\pi\right)^{2},&0\leq|\boldsymbol{p}|\leq k_{F},\\ 2R\left(k_{F}a\over\pi\right)^{2}\big{(}{k_{F}\over|\boldsymbol{p}}|\big{)}^{4},&k_{F}<|\boldsymbol{p}|\leq 4\,{\rm GeV},\\ 0,&|\boldsymbol{p}|>4\,{\rm GeV},\end{cases}$ (5)
where $a\equiv 2\,{\rm GeV}^{-1}$ is a characteristic size, $C\equiv{4\over 3}\pi k_{F}^{3}$ the phase space volume, and $R\equiv 1/(1-k_{F}/4\,{\rm GeV})$, respectively. The typical Fermi momentum for the Xenon nuclei is $k_{F}\approx 0.26\,\rm GeV$ which corresponds to $1/4$ of the speed of light. Nevertheless, the correction induced by the nucleon velocity $v_{N}$ to the cross section only appears at the quadratic order $\mathcal{O}(v^{2}_{N})$. Being below 10 %, it is safe to omit this correction for an estimation of the projected sensitivity at DM and neutrino experiments. Since the dimension-6 four-fermion operators in Eq. (1) scales as $1/\Lambda^{2}_{\rm eff}$, the cross section in Eq. (4) has $1/\Lambda^{4}_{\rm eff}$ scaling. A 10 % deviation in event rate can only lead to a negligible 2.5% uncertainty in the projected sensitivity reach of $\Lambda_{\rm eff}$. Being different from the fermionic absorption DM case Ge:2024euk , the nucleon consumption cross section is not affected by the nucleon Fermi motion or Pauli blocking effect since there is no nucleon in the final state.
Search at DM and Neutrino Experiments – As mentioned earlier, the proton consumption case will leave two visible particles, a charged lepton and a daughter nucleus in the excited state. Since both proton and the DM particles are heavy with $\mathcal{O}($GeV) masses, the released energy is mainly carried away by the light charged lepton. This also applies for the neutron consumption scenario. Then the charged leptons can directly deposit energy as Cherenkov light in a water detector such as Super-K or Hyper-K as well as scintillation light in others. In addition, positron $e^{+}$ can annihilate with a material electron to create two 511 keV gamma photons. For $\mu^{+}$, it may also decay inside the detector to produce a Michel positron first. Finally, the daughter nucleus in the excited state will emit a gamma photon. Altogether, the proton consumption can leave rich signal phenomena with double or even triple coincidence that can be used to uniquely select out the signal and at the same time significantly suppress backgrounds. It is then safe to assume background-free search for a first try of sensitivity estimation.
Probing the proton consumption can be achieved at DM direct detection experiments such as PandaX, XENON, and LZ with low energy thresholds at keV scale. On the other hand, the neutrino experiments like Super-K, Hyper-K, and JUNO have much larger size to compensate their higher thresholds. We take PandaX-4T (4 ton of 131Xe) PandaX:2018wtu , DUNE (40 kton of 40Ar) DUNE:2020lwj ; DUNE:2020ypp , JUNO (20 kton of $\rm C_{18}H_{30}$) JUNO:2015zny , and Super-K (22 kton of H2O) or Hyper-K (190 kton of H2O) Hyper-Kamiokande:2018ofw to demonstrate the projected sensitivities at Xenon, Argon, liquid scintillator, and water based experiments, respectively. The event rate is,
${d^{2}N\over dMdt}={N_{A}\times N_{T/p}\over\tilde{m}_{T}}\rho_{\chi}{\langle\sigma v_{\chi}\rangle\over m_{p}x},$ (6)
where the DM number density $\rho_{\chi}/m_{\chi}$ has been replaced by $\rho_{\chi}/m_{p}x$ which combines with the $f(x)$ function of $\sigma v_{\chi}$ to give $f(x)/x$ as mentioned earlier. Besides the Avogadro number $N_{A}=6.022\times 10^{23}$, $N_{T/p}$ is the proton number in the target $T$ while $\tilde{m}_{T}$ is the target molar mass,
$N_{\rm{}^{131}Xe/p}=54,$ $\quad\tilde{m}_{\rm{}^{131}Xe}=131.292\,{\rm g/mol},$ (7a)
$N_{\rm{}^{40}Ar/p}=18,$ $\quad\tilde{m}_{\rm{}^{40}Ar}=39.948\,{\rm g/mol},$ (7b)
$N_{\rm C_{18}H_{30}/p}=138,$ $\quad\tilde{m}_{\rm C_{18}H_{30}}=246.4308\,{\rm g/mol},$ (7c)
$N_{\rm H_{2}O/p}=10,$ $\quad\tilde{m}_{\rm H_{2}O}=18.02\,{\rm g/mol},$ (7d)
for the four experiments, respectively.
Figure 3: The solid (dashed) color curves show the experimental sensitivity on the effective scale $\Lambda_{\tt eff}$ from different experiments by assuming 1 year exposure with 100 (1) events. The upper (lower) plot is for the interactions with scalar (pseudo-scalar) DM current.
With the DM local density $\rho_{\chi}=0.4\,\rm GeV/cm^{3}$, the event rates are roughly,
${d^{2}N\over dMdt}=\left(9.2,7.5,9.3,6.8\right){10^{5}\over\rm ton\cdot yr}\left|1\,\rm TeV\over\Lambda_{\tt eff}\right|^{4}\frac{f^{i}(x)}{x},$ (8)
for the different target materials. Considering the huge volume of 2.25 kton and 4 ton for Super-K and PandaX-4T, the total event number predicted with $\Lambda_{\tt eff}=1$ TeV can reach $10^{10}$ and $10^{6}$ level within 1 year which is quite significant. This indicates that the projected sensitivity at DM and neutrino experiments can be much higher than 1 TeV.
Fig. 3 shows the projected sensitivity by requiring the signal event number needs to reach at least 100 (solid) or 1 (dashed). For a light sub-GeV DM, the effective scale can be probed up to $\sim 10$ TeV at the ton-scale PandaX-4T and the impressive $\gtrsim 500$ TeV at Hyper-K for 100 events. In the sub-GeV range, there is no big difference between the scalar ($\mathcal{O}^{\tt SS(P)}_{\chi\ell p}$) and pseudo-scalar ($\mathcal{O}^{\tt PS(P)}_{\chi\ell p}$) DM currents since the $x$-dependent function $f(x)/x$ does not differ much as shown in Fig. 2. With increasing DM mass, the probed scale of $\Lambda_{\tt eff}$ keeps decreasing since the $x$-dependent $f(x)/x$ and hence the event rate keeps decreasing. Between the scalar and pseudo-scalar DM current operators, the sensitivity deceases faster for the pseudo-scalar one, being consistent with the faster decreases of $f(x)/x$ shown in Fig. 2. The sensitivity for the vector, axial-vector, and tensor operators is similar as the scalar case with an ${\cal O}(1)$ rescaling factor.
For the UV model provided earlier, the effective scale $\Lambda_{\tt eff}^{-2}\sim y_{1}y_{2}y_{3,4}\mu_{S}\Lambda_{\tt QCD}^{3}/(m_{S_{1}}m_{S_{2}}m_{S_{3}})^{2}$ is determined by not just the new physics scale, such as the $S_{i}$ masses $m_{S_{i}}$, but also the QCD condensation scale $\Lambda_{\tt QCD}$. The projected sensitivities then convert to a constraint on the scalar masses, $m_{S_{i}}\lesssim y^{1/2}_{S_{i}}\mu^{1/6}_{S}\Lambda^{1/2}_{\tt QCD}(\Lambda^{\tt max}_{\tt eff})^{1/3}$. Although the QCD scale is just around $200$ MeV, the sensivity on $m_{S_{i}}$ needs not to be small since $\mu_{S}$ can be large.
Discussion and Conclusions – Not just the DM kinetic energy can induce nucleus recoil with elastic scattering, the particle mass can also be absorbed and then converted to the recoil energy. The previous studies mainly focus on the dark sector particles. Our study shows that consuming the nucleons inside a detector target can provide even more efficient conversion from mass to the energy deposit. For comparison, the fermionic absorption DM can only leave a recoil energy $T_{r}=m^{2}_{\chi}/2(m_{T}+m_{\chi})$ where $m_{T}$ is the target (electron, nucleon, or nucleus) mass. Since $m_{\chi}$ can not be too large ($m_{\chi}\ll m_{N}$) such that the fermionic absorption DM can be stable enough to survive until today, the recoil energy is actually much smaller than the DM mass, $T_{r}\ll m_{\chi}$. Since the nucleon consumption DM mass receives a lower bound, $m_{\chi}\gtrsim m_{p}/2$, to guarantee the proton stability, our new scenario of proton consumption can give the charged lepton an energy of $E_{\ell}=(m^{2}_{p}+2m_{p}m_{\chi}+m^{2}_{\ell})/2(m_{p}+m_{\chi})$ which needs not to be suppressed and the mass conversion to deposit energy is very efficient.
Notes Added – During the finalization of this paper, another paper Ema:2024wqr appeared on arXiv and briefly touched the similar idea of nucleon consumption DM. Although the title of their paper emphasized that baryon destruction is “catalyzed” by DM, the scenario of two-component DM has a major difference from ours. Our scheme provides a more economical realization with just a single DM component. The released mass purely comes from the consumed nucleon without resorting to the DM particle.
Acknowledgements
The authors would like to thank Yong Du for useful discussions and clarifying the effect of Fermi motion. SFG is supported by the National Natural Science Foundation of China (12375101, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604). SFG is also an affiliate member of Kavli IPMU, University of Tokyo. XDM is supported by the Guangdong Major Project of Basic and Applied Basic Research (Grant No.2020B0301030008) and by the NSFC (Grant No. NSFC-12305110).

### v2

@@
 Abstract
-We propose the nucleon consumption induced by dark matter as a new scenario to overcome the energy threshold of direct detection.
+We propose the nucleon consumption induced by dark matter (DM) as a new scenario to overcome the energy threshold of direct detection.
 It can be realized with proton ($\chi+p\rightarrow\chi+\ell^{+}$) or neutron ($\chi+n\rightarrow\chi+\nu$) target.
@@
 Especially, the weakly interacting massive particle (WIMP), whose stability is usually protected by discrete symmetry, is a theoretically well-motivated scenario that can meet the required properties to address the DM conundrum and at the same time have a detectable possibility in terrestrial experiments Roszkowski:2017nbc ; Schumann:2019eaa .
-However, the DM direct detection experiments have not found signal but constrain the DM-nucleus cross section to an unprecedented level for the WIMP mass from GeV to hundreds of GeV Billard:2021uyg ; Misiaszek:2023sxe .
+However, the DM direct detection experiments have not found the signal but constrain the DM-nucleus cross section to an unprecedented level for the WIMP mass from GeV to hundreds of GeV Billard:2021uyg ; Misiaszek:2023sxe .
 Since the DM-nucleus elastic scattering can only convert the DM kinetic energy to the nucleus recoil energy, it is difficult for a sub-GeV light DM to overcome the energy threshold of a direct detection experiment.
@@
 On the theoretical side, various DM scenarios have appeared.
-The recently proposed fermionic absorption DM is another example of light DM particle that can inject enough energy to the detector to produce a detectable signal with either nuclei Dror:2019onn ; Dror:2019dib ; Li:2022kca ; Ma:2024aoc ; Ge:2024euk ; Ma:2024tkt or electron Dror:2020czw ; Ge:2022ius ; Li:2022kca ; Ge:2023wye target.
+The detection threshold applies not just for sub-GeV DM but also for heavier DM whose events without large enough recoil energy cannot be detected.
+The recently proposed fermionic absorption DM is another example of a light DM particle that can inject enough energy to the detector to produce a detectable signal with either nuclei Dror:2019onn ; Dror:2019dib ; Li:2022kca ; Ma:2024aoc ; Ge:2024euk ; Ma:2024tkt or electron Dror:2020czw ; Ge:2022ius ; Li:2022kca ; Ge:2023wye target.
 The idea is that the DM ($\chi$) is absorbed by nuclear or atomic target ($T$) while emitting an undetectable neutrino ($\nu$), $\chi+T\to\nu+T^{\prime}$, while the target ($T^{\prime}$) recoils.
 Such signals have been searched at PandaX-4T PandaX:2022osq ; PandaX:2022ood , Majorana Majorana:2022gtu , EXO-200 EXO-200:2022adi , and CDEX CDEX:2022rxz ; CDEX:2024bum .
-Using the fermion absorption and converting its mass to recoil energy for detecting an unkown particle has a long history back to the 1940s.
-Although Pauli also pointed out the possibility of using its electromagnetic dipole interactions to probe the neutrino he proposed in the same letter Pauli:1991owm , both its very existence and electromagnetic interactions are hypothetical.
+Using the fermion absorption and converting its mass to recoil energy for detecting an unknown particle has a long history back to the 1940s.
+Pauli also pointed out the possibility of using its electromagnetic dipole interactions to probe the neutrino he proposed in the same paper Pauli:1991owm .
+However, both its very existence and electromagnetic interactions were hypothetical at that time.
 With only weak interactions, neutrinos are extremely difficult to detect.
-So Kan Chang Wang proposed in 1942 a very smart way of using the $K$-shell electron capture and detecting the mono-energetic daughter nuclei recoil to probe the existence of neutrino Wang:1942huv ; Wang:1947 .
+So Kan Chang Wang proposed in 1942 a very smart way of using the $K$-shell electron capture $e^{-}+{}^{Z}_{A}{\rm N}\rightarrow{}^{Z-1}_{~{}~{}A}{\rm N}+\nu$ and detecting the mono-energetic daughter nuclei recoil to probe the existence of neutrino Wang:1942huv ; Wang:1947 .
 This is essentially absorbing/consuming the known electron and releasing its mass as energy to probe the unknown neutrino.
@@
 Since one nucleon is consumed in the process to release almost 1 GeV energy, we anticipate there will be enough energy deposit in the detector and thus can be searched for at the DM direct detection and neutrino experiments.
-Nucleon Consumption Operators and Models – For the nucleon consumption scenarios induced by DM, it may happen at the nucleus level if the momentum transfer is not large enough.
-Then the target nucleus ($N$) reduces to a lighter one ($N^{\prime}$), $\chi+N\rightarrow\chi+N^{\prime}+\ell/\nu$ where the final-state charged lepton ($\ell$) or neutrino ($\nu$) is necessary to balance the fermion flow and angular momentum.
-However, it is not easy to make the momentum transfer transparent with a three-body final state.
+Nucleon Consumption – If the nucleon consumption happens at the nuclei level, the target nucleus ($N$) reduces to a lighter one ($N^{\prime}$), $\chi+N\rightarrow\chi+N^{\prime}+\ell/\nu$, where the final-state charged lepton ($\ell$) or neutrino ($\nu$) is necessary to balance the fermion flow and angular momentum.
 On the other hand, a large enough momentum transfer would render the nucleon consumption process to happen at the nucleon level.
@@
 For the DM mass below half of the nucleon mass, the same interaction that induces nucleon consumption can also lead to the proton ($p\rightarrow 2\chi+\ell^{+}$) or neutron ($n\to 2\chi+\nu$) decay.
-The subsequent dark decay of neutron can address the longstanding discrepancy between the bottle- and beam-type measurements of the neutron lifetime ParticleDataGroup:2022pth ; Yue:2013qrc .
-However, the nucleon decay can be forbidden by requiring a heavier DM mass to be on the safe side.
+To avoid complication, we impose a lower limit on the DM mass, $m_{\chi}\gtrsim(m_{p},m_{n})/2$, to guarantee stable nucleons.
 With a nonrelativistic incoming DM particle and the initial proton at rest, the consumed proton mass is released into the final-state DM and charged lepton energies $E_{\chi}=(m^{2}_{p}+2m_{p}m_{\chi}+2m^{2}_{\chi}-m^{2}_{\ell})/2(m_{p}+m_{\chi})$ and $E_{\ell}=(m^{2}_{p}+2m_{p}m_{\chi}+m^{2}_{\ell})/2(m_{p}+m_{\chi})$, respectively.
@@
 So we would focus on the proton consumption scenario for illustration while most results also apply for the neutron case.
-The nucleon consumption can be described in terms of four-fermion effective operators involving a pair of fermion DM particles, proton, and charged lepton are,
+The nucleon consumption can be described in terms of four-fermion effective operators involving a pair of fermion DM particles, proton, and charged lepton,
 ${\cal O}_{\chi\ell p}^{\tt SS(P)}$ $=\overline{\chi}\chi\,\overline{\ell^{c}}(i\gamma_{5})p,$ (1a)
@@
 where $\ell^{c}$ is the charge conjugation of the lepton field.
-For the superscript, the first character stands for the Lorentz type of the DM current while the second and the one in parenthesis for the proton-lepton current.
+For the superscript, the first character stands for the Lorentz type of the DM current while the second and the one in parentheses for the proton-lepton current.
 The DM current can also take the fermion number violating case, $\overline{\chi^{c}}\Gamma\chi$.
 Then only the scalar, pseudo-scalar, and axial-vector operators ${\cal O}_{\chi\ell p}^{\tt S,P,A}$ can survive.
+In such a case, additional discrete symmetries can be imposed to forbid unwanted processes such as single nucleon decay, which will be realized in the toy model given below.
 The strength of each operator is parametrized by a Wilson coefficient $C_{\chi\ell p}^{i}$.
 For the DM-neutron consumption, one may just replace the charged lepton-proton current by the neutrino-neutron current, $(\overline{\nu}\Gamma n)$ or $\overline{\nu^{c}}\Gamma n$.
-The nucleon consumption operators in Eq.
-(1) can be realized by extending the SM with a Dirac fermion DM candidate $\chi(1,1,0)$ besides three leptoquarks $S_{1}(3,1,2/3)$, $S_{2}(3,1,-1/3)$, and $S_{3}(3,1,-1/3)$.
-In addition, one needs to impose a discrete $\mathbb{Z}_{4}$ symmetry under which the charge assignments are,
-$\chi:i,\quad S_{1,2}:-i,\quad S_{3},L,e:-1,$ (2)
-while the remaining SM particles are $\mathbb{Z}_{4}$ singlets. Here $L$ and $e$ are the SM left- and right-handed leptons. The relevant new physics Lagrangian is
-${\cal L}_{\tt NP}$ $\ni y_{1}(\overline{u}\chi_{L})S_{1}+y_{2}(\overline{d}\chi_{L})S_{2}+y_{3}(\overline{Q}\epsilon L^{c})S_{3}$
-$+y_{4}(\overline{u}e^{c})S_{3}-\mu_{S}\epsilon_{ijk}S_{1}^{i}S_{2}^{j}S_{3}^{k}+{\rm H.c.},$ (3)
-where $Q,u,d$ are the SM quarks while $i,j,k$ are the color indices in the fundamental representation of $\rm SU(3)_{c}$.
-For the Yukawa couples $y_{1,2,3,4}$, the flavor indices have been omitted for simplicity.
-The dimensional $\mu_{S}$ is the triple scalar coupling.
-[…]
-Figure 1: The Feynman diagram for generating the effective operators responsible for the nucleon consumption induced by DM.
-The Feynman diagram responsible for the nucleon comsuming DM is shown in Fig.
-1, which will generate operators with (pseudo-)scalar DM currents in Eq.
-(1) once quarks condensate into nucleons.
-While the couplings with $S_{1}$ and $S_{2}$ already contribute one up and one down quarks, whether the model induces a proton or neutron consumption is determined by the $S_{3}$ Yukawa couplings.
-A vanishing $y_{3}$ prevents neutrino from participating the process and hence only leads to the proton consumption scenario with nonzero $y_{4}$.
-On the other hand, a nonzero $y_{3}$ can induce both proton and neutron consumption.
-Cancellation between the $y_{3}$ and $y_{4}$ contributions to the proton decay can also happen by tuning values.
-However, it is more desirable to adjust the model with a light sterile neutrino $\nu_{R}(1,1,0)$ that has a $\mathbb{Z}_{4}$ charge $-1$.
-Furthermore, the SM leptons and quarks all are neutral under the $\mathbb{Z}_{4}$ symmetry.
-In this case, the Yukawa terms associated with $y_{3,4}$ are replaced by $y_{3}(\overline{d}\nu_{R}^{c})S_{3}$.
 Consumption at Rest and Fermi Motion – Since the halo DM particles travel at only $10^{-3}$ of the speed of light, it is a good approximation to treat them as at rest.
@@
 As shown in Fig.
-2, the $x$-dependent $f(x)/x$ that finally appears in the event rate is typically in the range of $10^{-3}\sim 10$.
+1, the $x$-dependent $f(x)/x$ that finally appears in the event rate is typically in the range of $10^{-3}\sim 10$.
 The size of $f(x)/x$ is sizable with the only exception of the pseudo-scalar case.
@@
 Being different from the fermionic absorption DM case Ge:2024euk , the nucleon consumption cross section is not affected by the nucleon Fermi motion or Pauli blocking effect since there is no nucleon in the final state.
-Search at DM and Neutrino Experiments – As mentioned earlier, the proton consumption case will leave two visible particles, a charged lepton and a daughter nucleus in the excited state.
-Since both proton and the DM particles are heavy with $\mathcal{O}($GeV) masses, the released energy is mainly carried away by the light charged lepton.
-This also applies for the neutron consumption scenario.
-Then the charged leptons can directly deposit energy as Cherenkov light in a water detector such as Super-K or Hyper-K as well as scintillation light in others.
-In addition, positron $e^{+}$ can annihilate with a material electron to create two 511 keV gamma photons.
-For $\mu^{+}$, it may also decay inside the detector to produce a Michel positron first.
-Finally, the daughter nucleus in the excited state will emit a gamma photon.
-Altogether, the proton consumption can leave rich signal phenomena with double or even triple coincidence that can be used to uniquely select out the signal and at the same time significantly suppress backgrounds.
-It is then safe to assume background-free search for a first try of sensitivity estimation.
+[…]
 Probing the proton consumption can be achieved at DM direct detection experiments such as PandaX, XENON, and LZ with low energy thresholds at keV scale.
@@
 2 shows the projected sensitivity by requiring the signal event number needs to reach at least 100 (solid) or 1 (dashed).
-For a light sub-GeV DM, the effective scale can be probed up to $\sim 10$ TeV at the ton-scale PandaX-4T and the impressive $\gtrsim 500$ TeV at Hyper-K for 100 events.
+For a light sub-GeV DM, the effective scale can be probed up to $\sim 10$ TeV at the ton-scale PandaX-4T and the impressive $\gtrsim 200$ TeV at Hyper-K for 100 events.
 In the sub-GeV range, there is no big difference between the scalar ($\mathcal{O}^{\tt SS(P)}_{\chi\ell p}$) and pseudo-scalar ($\mathcal{O}^{\tt PS(P)}_{\chi\ell p}$) DM currents since the $x$-dependent function $f(x)/x$ does not differ much as shown in Fig.
-2.
+1.
 With increasing DM mass, the probed scale of $\Lambda_{\tt eff}$ keeps decreasing since the $x$-dependent $f(x)/x$ and hence the event rate keeps decreasing.
 Between the scalar and pseudo-scalar DM current operators, the sensitivity deceases faster for the pseudo-scalar one, being consistent with the faster decreases of $f(x)/x$ shown in Fig.
-2.
-The sensitivity for the vector, axial-vector, and tensor operators is similar as the scalar case with an ${\cal O}(1)$ rescaling factor.
-For the UV model provided earlier, the effective scale $\Lambda_{\tt eff}^{-2}\sim y_{1}y_{2}y_{3,4}\mu_{S}\Lambda_{\tt QCD}^{3}/(m_{S_{1}}m_{S_{2}}m_{S_{3}})^{2}$ is determined by not just the new physics scale, such as the $S_{i}$ masses $m_{S_{i}}$, but also the QCD condensation scale $\Lambda_{\tt QCD}$.
-The projected sensitivities then convert to a constraint on the scalar masses, $m_{S_{i}}\lesssim y^{1/2}_{S_{i}}\mu^{1/6}_{S}\Lambda^{1/2}_{\tt QCD}(\Lambda^{\tt max}_{\tt eff})^{1/3}$.
-Although the QCD scale is just around $200$ MeV, the sensivity on $m_{S_{i}}$ needs not to be small since $\mu_{S}$ can be large.
+1.
+The sensitivity for the vector, axial-vector, and tensor operators is similar to the scalar case with an ${\cal O}(1)$ rescaling factor.
+Toy Model – We use the leptoquark model Dorsner:2016wpm to provide a concrete realization of the nucluon consumption operators in Eq.
+(1).
+Besides a Dirac fermion DM candidate $\chi(1,1,0)$, the SM is extended with three leptoquarks $S_{1}(3,1,2/3)$, $S_{2}(3,1,-1/3)$, and $S_{3}(3,1,-1/3)$, where the three numbers in parentheses are their charges under the SM gauge groups $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$.
+In addition, one needs to impose a discrete $\mathbb{Z}_{4}$ symmetry under which the charge assignments are,
+$\chi:-i,\quad S_{1,2}:-i,\quad S_{3},L,e:-1,$ (7)
+while the remaining SM particles are $\mathbb{Z}_{4}$ singlets. Here $L$ and $e$ are the SM left- and right-handed leptons. The relevant new physics Lagrangian is
+${\cal L}_{\tt NP}$ $\ni y_{1}(\overline{u}\chi_{R}^{c})S_{1}+y_{2}(\overline{d}\chi_{R}^{c})S_{2}+y_{3}(\overline{Q}\epsilon L^{c})S_{3}$
+$+y_{4}(\overline{u}e^{c})S_{3}-\mu_{S}\epsilon_{ijk}S_{1}^{i}S_{2}^{j}S_{3}^{k}+{\rm H.c.},$ (8)
+where $Q,u,d$ are the SM quarks while $i,j,k$ are the color indices in the fundamental representation of $\rm SU(3)_{c}$.
+For the Yukawa couples $y_{1,2,3,4}$, the flavor indices have been omitted for simplicity.
+The dimensional triple scalar coupling $\mu_{S}$ is largely a free parameter.
+[…]
+Figure 3: The Feynman diagram for generating the effective operators responsible for the nucleon consumption induced by DM.
+The Feynman diagram responsible for the nucleon consuming DM is shown in Fig.
+3, which will generate operators with (pseudo-)scalar DM currents in Eq.
+(1) once quarks condensate into nucleons.
+While the couplings with $S_{1}$ and $S_{2}$ already contribute one up and one down quarks, whether the model induces a proton or neutron consumption is determined by the $S_{3}$ Yukawa couplings.
+A vanishing $y_{3}$ prevents neutrino from participating in the process and hence only leads to the proton consumption scenario with nonzero $y_{4}$.
+On the other hand, a nonzero $y_{3}$ can induce both proton and neutron consumption.
+Cancellation between the $y_{3}$ and $y_{4}$ contributions to the proton consumption can also happen by tuning values.
+However, it is more desirable to adjust the model with a light sterile neutrino $\nu_{R}(1,1,0)$ that has a $\mathbb{Z}_{4}$ charge $-1$.
+Furthermore, the SM leptons and quarks all are neutral under the $\mathbb{Z}_{4}$ symmetry.
+In this case, the Yukawa terms associated with $y_{3,4}$ are replaced by $y_{3}(\overline{d}\nu_{R}^{c})S_{3}$.
+The cut-off scale $\Lambda_{\tt eff}^{-2}\sim y_{1}y_{2}y_{3,4}\mu_{S}\Lambda_{\tt QCD}^{3}/(m_{1}m_{2}m_{3})^{2}$ is determined by not just the new physics scale, such as the $S_{i}$ masses $m_{i}$, but also the QCD condensation scale $\Lambda_{\tt QCD}$.
+The projected sensitivities then convert to a constraint on the scalar masses, $m_{i}\lesssim y^{1/2}_{S_{i}}\mu^{1/6}_{S}\Lambda^{1/2}_{\tt QCD}(\Lambda^{\tt max}_{\tt eff})^{1/3}$.
+Although the QCD scale is just around $200$ MeV, the sensitivity on $m_{i}$ needs not to be small since $\mu_{S}$ can be large.
+Currently, the collider bounds on the leptoquark masses $m_{i}$ are around 2 TeV depending on the decay assumption of leptoquarks ATLAS:2020dsk ; ATLAS:2024yxs .
+Such constraints translate into a bound on $\Lambda_{\rm eff}\sim 300\,\rm TeV$ for ${\cal O}(1)$ Yukawa couplings and ${\cal O}(10^{11}\,\rm GeV)$ of $\mu_{S}$.
 Discussion and Conclusions – Not just the DM kinetic energy can induce nucleus recoil with elastic scattering, the particle mass can also be absorbed and then converted to the recoil energy.
@@
 Our study shows that consuming the nucleons inside a detector target can provide even more efficient conversion from mass to the energy deposit.
+This can overcome not just the $\mathcal{O}$(keV) energy threshold at the DM experiments but also even the $\mathcal{O}$(MeV) one at neutrino experiments.
 For comparison, the fermionic absorption DM can only leave a recoil energy $T_{r}=m^{2}_{\chi}/2(m_{T}+m_{\chi})$ where $m_{T}$ is the target (electron, nucleon, or nucleus) mass.
@@
 Notes Added – During the finalization of this paper, another paper Ema:2024wqr appeared on arXiv and briefly touched the similar idea of nucleon consumption DM.
-Although the title of their paper emphasized that baryon destruction is “catalyzed” by DM, the scenario of two-component DM has a major difference from ours.
+Although the title of their paper emphasized that baryon destruction is catalyzed by DM, the scenario of two-component DM has a major difference from ours.
 Our scheme provides a more economical realization with just a single DM component.
 The released mass purely comes from the consumed nucleon without resorting to the DM particle.
+Although their model can lead to nucleon consumption with dinucleon initial state, $n+n\to n+\nu$, such reaction together with the single nucleon decays like $p\to\pi^{0}e^{+}$ are forbidden in our case due to the exact $\mathbb{Z}_{4}$ symmetry.
+The constraint obtained from the DM capture in neutron star and the subsequent heating due to nucleon consumption can also be relaxed.
+For the neutron consumption case, the DM capture in neutron star with elastic scattering can be induced by the mediation of both $S_{1}$ and $S_{2}$ where cancellation can happen to make the capture rate small enough.
+The neutron star heating effect can be further minimized if the Yukawa terms associated with $y_{3,4}$ in Eq.
+(8) are replaced by $y_{3}(\bar{d}\nu_{R}^{c})S_{3}$ since the right-handed neutrino is sterile and hence cannot deposit energy.
+For the proton consumption case, with ${\cal O}(1\sim 10)\%$ suppression in the proton number density around the neutron star surface, the heating constraint is less stringent.
+In addition, the captured DM needs not to stay around the neutron star surface and can diffuse deep into the neutron star body, which would lead to a further suppression.
+Since the sizable fraction of proton density only appears in a very narrow layer around the neutron star surface, such suppression due to DM diffusion is quite significant.
 Acknowledgements
 The authors would like to thank Yong Du for useful discussions and clarifying the effect of Fermi motion.
-SFG is supported by the National Natural Science Foundation of China (12375101, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604).
+SFG is supported by the National Natural Science Foundation of China (12375101, 12425506, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604).
 SFG is also an affiliate member of Kavli IPMU, University of Tokyo.